OpenStudy (anonymous):

How do I solve 16 int_{0}^{pi/6} sqrt{1-sin ^2 theta} cos theta d theta ?

OpenStudy (anonymous):

\[16 \int\limits_{0}^{\pi/6} \sqrt{1-\sin ^2 \theta} \cos \theta d \theta\]

OpenStudy (anonymous):

@.Sam. @UnkleRhaukus

OpenStudy (anonymous):

I was going to let sin theta = u, but then I got stuck.

OpenStudy (phi):

change the radical to cos theta you get cos^2 theta d theta now use the identity cos^2(x) = (1+ cos(2x))/2

OpenStudy (anonymous):

ohhh...

OpenStudy (phi):

cos^2 x + sin^2 x = 1

OpenStudy (anonymous):

Just a sec.

OpenStudy (anonymous):

okay, now I have \[8 \int\limits_{0}^{\pi/6} 1+ \cos^2 \theta d \theta \]

OpenStudy (phi):

yes

OpenStudy (anonymous):

Wait, that should be cos(2 theta)

OpenStudy (phi):

yes 2 theta

OpenStudy (anonymous):

Now I have \[8[\theta - (1/2) \sin(2 \theta) from 0 \to \pi/6\]

OpenStudy (phi):

ok

OpenStudy (phi):

though I think that should be + sin(2x)

OpenStudy (phi):

as a check: d (+sin 2x) = cos 2x 2 dx